In a population of 225 women, the heights of the women are normally distributed with a mean of 64.5 inches and a standard deviation of 2.9 inches. If 25 women are selected at a random, find the probability that their mean height will exceed 66 inches. Assume that the sampling is done without replacement and use a finite population correction factor with N=225. Pls. show solution
Answer #1

N = 225

n = 25

mu=64.5

sigma= 2.9

Finite Population Correction Factor (FPC) is got as follows:

SE=rac{sigma }{sqrt{n}}sqrt{rac{N-n}{N-1}}

Substituting values, we get:

尚1222-25 =0.58 SE2.g V25 V 22510.58,/200 = 0.5480

To find P(ar{x}>66):
Z = (66 - 64.5)/0.5480 = 2.7370

Table of Area Under Standard Normal Curve gives area = 0.4969

So,

P(ar{x} > 66) =0.5 - 0.4969 = 0.0031

So,

Answer is:

0.0031